OpenStudy (anonymous):

system of linear system: y=-x+15 y=x+1

OpenStudy (anonymous):



OpenStudy (phi):

if -x+15 equals y and x+1 equals y then -x+15 = x+1 can you solve this for x?

OpenStudy (anonymous):

would you add x to both sides

OpenStudy (phi):

that is a good step: -x +15 +x = x+1 +x

OpenStudy (phi):

what do you get after simplifying?

OpenStudy (anonymous):

i really need to be taught step by step

OpenStudy (phi):

first do the left side of -x +15 +x = x+1 +x if instead of x we had 2: -2 + 15 +2 what is that?

OpenStudy (anonymous):

ratio 2:-2

OpenStudy (phi):

no, it is just adding (but using negative numbers) -2 + 15 +2 is 15 (-2 + 2 is 0 and 0+15 is 15) it is the same for x: -x + 15 + x is the same as x + -x +15, x + -x is 0 and 0+15 is 15 -x +15 +x = x+1 +x becomes 15 = x+1+x we can write this as 15= x+x+1 (just switched the order) do you know what x+x is?

OpenStudy (anonymous):

x^2

OpenStudy (phi):

x^2 means x*x (x times x) this is x plus x if you have 1 x and add one more x how many x's do you have?

OpenStudy (anonymous):

2x

OpenStudy (phi):

yes so now we have 15= x+x+1 becomes 15 = 2x + 1 what happens if you add -1 to both sides?

OpenStudy (anonymous):

but were suppose to get an ordered pair

OpenStudy (phi):

first things first. 15 = 2x + 1 what happens if you add -1 to both sides?

OpenStudy (anonymous):

i got 14/2 which equals 7

OpenStudy (phi):

15 -1 = 2x +1 -1 15-1 is 14 14= 2x + 1 -1 1-1 is 0 14= 2x+0 2x+0 is 2x 14= 2x divide both sides by 2 14/2 = 2x/2 7= x yes, that is good. now you need y

OpenStudy (anonymous):

so i would subsitute x for 7 right

OpenStudy (phi):

yes. you can use either equation (they both better give the same answer for y) I would use y=x+1 because it looks easier to figure out

OpenStudy (anonymous):

so it would be 7 and 7

OpenStudy (phi):

how?

OpenStudy (phi):

y=x+1

OpenStudy (anonymous):

so i plugged in the 7 for x and got 8 sorry

OpenStudy (phi):

yes, much better x=7, y=8 is the answer.

OpenStudy (phi):

or (7,8) if they want an "ordered pair"

OpenStudy (anonymous):

would you mind helping me with two more please @phi

OpenStudy (phi):

please make it a new post